OpenStudy (anonymous):

Solve the equation for y. 7x - 3y = 4 A. y = x + B. y = 7x - 4 C. y = x + D. y = x -

OpenStudy (anonymous):

7x - 3y = 4 -3y=4-7x y=(4-7x)/-3

OpenStudy (anonymous):